# Foundation Models for Autonomous Driving System: An Initial Roadmap

### v1

@@ 2.1. High-quality Dataset for Autonomous Driving
Challenge I: Dataset Cleaning and Curation. Dataset cleaning and curation serve as a critical foundation for developing FMs, ensuring data integrity, privacy protection, and efficient training. Key challenges include protecting privacy (Zhu et al., 2024a; Zhang et al., 2024b; Carlini et al., 2021), and mitigating bias in training datasets (Llorca et al., 2024; Li et al., 2024a; Hort et al., 2024a). The privacy challenge involves both preventing personal data from appearing in training datasets and ensuring it cannot be inferred from model outputs (Kim et al., 2023; Zhu et al., 2024a). Bias in training data can lead models to perpetuate systemic inequities and may introduce safety risks for unrepresented groups when deployed in real-world scenarios (Llorca et al., 2024). These challenges present several key opportunities for further research.
@@
  • Opportunity: Bias Mitigation. Dataset bias poses a critical challenge for autonomous driving. Recent studies (Llorca et al., 2024) have revealed limited diversity in key demographic attributes (i.e., age, sex, and skin tone) within existing AD datasets. This lack of representation could lead to safety risks when deploying FMs trained on such datasets, particularly for underrepresented groups. While recent research has made progress in addressing these concerns (Li et al., 2024a; Hort et al., 2024a, b; Chen et al., 2025), the detection and mitigation of dataset bias remains an important open challenge in autonomous driving.
@@
  • Opportunity: Privacy Preservation. Another critical research opportunity lies in developing effective privacy-preserving algorithms for FMs in autonomous driving. Established techniques such as differential privacy (Duan et al., 2023; Chua et al., 2024; Tang et al., 2024a; Tramèr et al., 2024), data cleaning (Kandpal et al., 2022; Carlini et al., 2021), and federated learning (Xu et al., 2024a; Zhang et al., 2024b; Sun et al., 2024b), have advanced the field, yet they consistently face challenges in balancing privacy preservation with data utility. The challenge is particularly acute with FMs, which tend to memorize training data extensively, potentially leading to privacy leakage even with the data used in fine-tuning processes (Zhu et al., 2024a). There is an urgent need for novel techniques that can ensure robust privacy guarantees while maintaining the comprehensive nature of training datasets required for large FMs.
@@
Challenge II: Augmenting Autonomous Driving Datasets. Despite significant investments in the development of autonomous driving datasets, current limitations in data quality and scale hinder their ability to comprehensively address the field’s challenges (Cui et al., 2024c). Moreover, certain critical scenarios remain difficult or nearly impossible to capture in real-world data collection (Wei et al., 2024c; Song et al., 2024a). These include high-risk situations such as accident aftermath and pedestrian-involved incidents. However, comprehensive testing of autonomous vehicles against these scenarios is crucial for safety validation. To overcome these challenges, researchers should increasingly explore methods for generating customized driving scenarios, enabling the effective simulation of these critical cases to augment existing datasets and enhance their utility.
@@
  • Opportunity: Data Management Framework for FMs. As FMs demonstrate performance improvements through data scaling and the significance of data becomes evident (Chowdhery et al., 2023), effective data management becomes increasingly critical. While various tools and methods have been proposed to explore how to properly manage the training data, encompassing data deduplication (Lee et al., 2022a), training data selection (Rae et al., 2022; Wang et al., 2024a; Li et al., 2024b), sampling high-quality data (Peng et al., 2025; Wettig et al., 2024), and dataset license compliance (Kim et al., 2024b; Jahanshahi and Mockus, 2025), there is still a lack of a unified framework and criteria. Although there have been some initial attempts in this area (Wang et al., 2024c; Ostendorff et al., 2024; Peng et al., 2025), systematic approaches to data management for FMs remain in their early stages. Given the massive scale and diverse sources of data required for training FMs, developing a comprehensive data management framework has become an urgent priority.
@@ 2.2. Computational Resources
Due to the computational-intensive nature of FMs, computational resources, including graphic processing units (GPUs), tensor processing units (TPUs), and other specialized AI accelerators, form the very foundation of the FM infrastructure. Building upon this hardware layer, distributed training frameworks and cloud computing enable efficient resource utilization and management. However, the complexity of distributed, computation-intensive training, and reliable efficient deployment introduces unique challenges and opportunities in adopting FMs in autonomous driving. In this section, we mainly discuss the challenges and opportunities related to hardware, issues with the software layer (e.g., distributed training framework) are discussed in Section 2.3.
Challenge & Opportunity: Trustworthy Hardware Design. The computational-intensive nature of FM training and inference necessitates reliance on proprietary and specialized hardware solutions. However, this dependency complicates ensuring compatibility, optimization, and security across the hardware stack (Wang et al., 2024b). The proprietary nature of hardware design and firmware creates a ”black-box” environment, where potential vulnerabilities may remain undetected (Yang et al., 2016; Wang et al., 2024b). Furthermore, the parallel processing architecture and shared resources make GPUs particularly vulnerable to hardware attacks (Jiang et al., 2017), potentially resulting in sensitive information leakage (e.g., model parameters (Naghibijouybari et al., 2021)) or even enabling arbitrary code execution (Lee et al., 2022b). Given that security is crucial for autonomous driving systems, there is an urgent need to design secure hardware that incorporates security features at the hardware level and can effectively resist hardware-level threats such as side-channel attacks.
@@ 2.3. Development and Deployment of FM4AD
Challenge & Opportunity: Understanding FM Development Toolchain. The development toolchain for FMs presents unprecedented complexity compared to traditional deep learning frameworks. The enormous scale of these models amplifies the intricacy of data pre-processing pipelines and distributed training systems (Wang et al., 2024b). Additionally, the fast-evolving nature of FM development hinders developers and researchers from maintaining a comprehensive understanding of the continuously evolving toolchain. A promising research avenue lies in mining the FM development toolchain. This approach entails a thorough examination and assessment of the libraries and tools employed in FM development. By analyzing the toolchain, developers can uncover inefficiencies and streamline processes while preserving functionality (Wang et al., 2024b; Nguyen et al., 2019; Liu et al., 2023a). Furthermore, this exploration may reveal gaps, fostering the design of innovative tools or enhancements to existing ones, better addressing the growing demands of FM development.
Challenge & Opportunity: Efficient and Reliable FM Deployment in Vehicle. With growing concerns regarding privacy and strict response time requirements for autonomous driving tasks, efficiently deploying FMs in vehicles has become increasingly important. Existing approaches have investigated model compression techniques such as pruning (Ma et al., 2023; Sun et al., 2024a), knowledge distillation (Gu et al., 2024; Kim et al., 2024a), and model quantization (Lin et al., 2024; Xi et al., 2025; Li et al., 2025), alongside inference optimization techniques including parallel computation (Shoeybi et al., 2020; Yu et al., 2024), KV cache (Kwon et al., 2023; Luohe et al., 2024), and request scheduling (Han et al., 2022; Yang et al., 2024b). While these techniques can alleviate computational burden and improve response times, they may introduce new vulnerabilities. For instance, researchers have identified tailored attacks targeting quantized models (Zhang et al., 2025; Yang et al., 2024a) and KV cache-based optimizations (Song et al., 2025; Wu et al., 2025). Consequently, there is an urgent need to develop techniques that ensure both safety and efficiency when deploying FMs in vehicles.
Challenge & Opportunity: Edge/Cloud Collaboration for FM Services. While quantization techniques can alleviate the computational burden of edge devices, model capability remains limited by available edge computing power. Conversely, cloud computing infrastructure offers high-performance processing for complex tasks but often struggles to meet privacy, reliability, and latency requirements crucial for autonomous driving. Recent research efforts have been focused on investigating collaboration between edge and cloud (Yang et al., 2024c; Hao et al., 2024; Zhang et al., 2024a; Tian et al., 2024b). By developing frameworks that intelligently schedule and coordinate tasks between edge and cloud resources, autonomous driving systems can achieve collaborative intelligence, enhance adaptability to varying conditions, leverage edge knowledge while preserving privacy, and optimize resource utilization across the edge-cloud continuum. Potential research directions include optimizing heterogeneous architecture fusion strategies, designing asynchronous update mechanisms, and robust communication schemes.

### v2

# Foundation Models for Autonomous Driving Systems: An Initial Roadmap

@@ 3.1. High-quality Dataset for Autonomous Driving
Challenge I: Dataset Cleaning and Curation. Dataset cleaning and curation serve as a critical foundation for developing FMs, ensuring data integrity, privacy protection, and efficient training. Key challenges include protecting privacy (Zhu et al., 2024b; Zhang et al., 2024b; Carlini et al., 2021), and mitigating bias in training datasets (Llorca et al., 2024; Li et al., 2024b; Hort et al., 2024a). The privacy challenge involves both preventing personal data from appearing in training datasets and ensuring it cannot be inferred from model outputs (Kim et al., 2023; Zhu et al., 2024b). Bias in training data can lead models to perpetuate systemic inequities and may introduce safety risks for unrepresented groups when deployed in real-world scenarios (Llorca et al., 2024). These challenges present several key opportunities for further research.
@@
  • Opportunity: Bias Mitigation. Dataset bias poses a critical challenge for autonomous driving. Recent studies (Llorca et al., 2024) have revealed limited diversity in key demographic attributes (i.e., age, sex, and skin tone) within existing AD datasets. This lack of representation could lead to safety risks when deploying FMs trained on such datasets, particularly for underrepresented groups. While recent research has made progress in addressing these concerns (Li et al., 2024b; Hort et al., 2024a, b; Chen et al., 2025a), these previous works mainly focus on the perception module (i.e., pedestrian detection), using fairness metrics such as Statistical Parity Differences (SPD), Equal Opportunity Difference (EOD), Average Odds Difference (AOD), and Miss Rate (MR). One open software engineering challenge lies in verifying how such biases propagate through the system pipeline. Future research could develop stratified, scenario-based testing suites where only demographic attributes vary, and evaluate whether downstream behaviors (e.g., safety margin, rule violations, intervention rate) exhibit systematic disparities, enabling fairness regression testing across releases. In addition, LLMs are known to exhibit bias and fairness issues (Gallegos et al., 2024); integrating FMs into ADS reasoning/planning may introduce new bias pathways beyond perception, which calls for dedicated test suites and systematic investigation.
@@
  • Opportunity: Privacy Preservation. Another critical research opportunity lies in developing effective privacy-preserving algorithms for FMs in autonomous driving. According to Staab et al. (Staab et al., 2024), LLMs can even infer personal attributes from real-world data, even when the text is anonymized using commercial tools. Wang et al. (Wang and Xu, 2026) highlight that malicious attackers can infer and trace vehicle trajectories in accident warning systems, raising serious privacy concerns and introducing potential security risks. Established techniques such as differential privacy (Duan et al., 2023; Chua et al., 2024; Tang et al., 2024a; Tramèr et al., 2024), data cleaning (Kandpal et al., 2022; Carlini et al., 2021), and federated learning (Xu et al., 2024a; Zhang et al., 2024b; Sun et al., 2024b), have advanced the field, yet they consistently face challenges in balancing privacy preservation with data utility. The challenge is particularly acute with FMs, which tend to memorize training data extensively, potentially leading to privacy leakage even with the data used in fine-tuning processes (Zhu et al., 2024b). There is an urgent need for novel techniques that can ensure robust privacy guarantees while maintaining the comprehensive nature of training datasets required for large FMs.
  • Opportunity: Machine Unlearning for FMs. While curating training data can mitigate issues upfront, FMs already deployed may contain private information or learned biases. For these models, machine unlearning emerges as a critical capability (Xu et al., 2023). The goal is to efficiently remove the influence of specific data points or concepts from a trained model without the need for a complete, costly retraining from scratch. This is essential for complying with data privacy regulations like the ”right to be forgotten” and for rectifying harmful biases discovered post-deployment (Xu et al., 2023). For coding tasks, machine unlearning has proven its effectiveness (Chu et al., 2025). The challenge of this opportunity lies in developing unlearning techniques that are not only effective at removing information but are also computationally efficient and maintain the model’s overall performance on other tasks, which is particularly difficult for FMs in autonomous driving, since these models inherently need to deal with multiple tasks.
@@
Challenge II: Augmenting Autonomous Driving Datasets. Despite significant investments in the development of autonomous driving datasets, current limitations in data quality and scale hinder their ability to comprehensively address the field’s challenges (Cui et al., 2024c). Moreover, certain critical scenarios remain difficult or nearly impossible to capture in real-world data collection (Wei et al., 2024c; Song et al., 2024a). These include high-risk situations such as accident aftermath and pedestrian-involved incidents. However, comprehensive testing of autonomous vehicles against these scenarios is crucial for safety validation. To overcome these challenges, researchers should increasingly explore methods for generating customized driving scenarios or automated data collection, enabling the effective simulation of these critical cases to augment existing datasets and enhance their utility.
@@
  • Opportunity: Data Management Framework for FMs. As FMs demonstrate performance improvements through data scaling and the significance of data becomes evident (Chowdhery et al., 2023), effective data management becomes increasingly critical. While various tools and methods have been proposed to explore how to properly manage the training data, encompassing data deduplication (Lee et al., 2022a), training data selection (Rae et al., 2022; Wang et al., 2024a; Li et al., 2024d), sampling high-quality data (Peng et al., 2025; Wettig et al., 2024), and dataset license compliance (Kim et al., 2024b; Jahanshahi and Mockus, 2025), there is still a lack of a unified framework and criteria. Although there have been some initial attempts in this area (Wang et al., 2024c; Ostendorff et al., 2024; Peng et al., 2025), systematic approaches to data management for FMs remain in their early stages. Given the massive scale and diverse sources of data required for training FMs, developing a comprehensive data management framework has become an urgent priority.
@@ 3.2. Computational Resources
Due to the computational-intensive nature of FMs, computational resources, including graphics processing units (GPUs), tensor processing units (TPUs), and other specialized AI accelerators, form the very foundation of the FM infrastructure. Building upon this hardware layer, distributed training frameworks and cloud computing enable efficient resource utilization and management. However, the complexity of distributed, computation-intensive training, and reliable efficient deployment introduces unique challenges and opportunities in adopting FMs in autonomous driving. In this section, we mainly discuss the challenges and opportunities related to hardware, issues with the software layer (e.g., distributed training framework) are discussed in Section 3.3.
Challenge I: Trustworthy Hardware Design. The computationally intensive nature of FMs necessitates a reliance on proprietary and specialized hardware, such as GPUs and other accelerators. This dependency creates significant challenges in ensuring security across the hardware stack, as the proprietary design of chips and firmware often results in a ”black-box” environment where vulnerabilities can remain undetected (Wang et al., 2024b; Yang et al., 2016). Furthermore, the parallel processing architectures and shared resources inherent to modern accelerators make them particularly susceptible to hardware-level attacks (Jiang et al., 2017). These attacks can lead to severe consequences, including the leakage of sensitive information like model parameters via side-channels (Naghibijouybari et al., 2021) or even enabling arbitrary code execution (Lee et al., 2022b). Given the paramount importance of security in autonomous driving, these hardware vulnerabilities present a fundamental risk that necessitates urgent research and development. Beyond the general exposure of proprietary accelerators, FM-centric ADS deployments introduce additional constraints that exacerbate the hardware trust problem. First, FM inference (and especially VLM/world-model pipelines) typically requires large model footprints (weights, intermediate activations, and KV caches), which can exceed the secure memory/enclave capacity of many practical TEE designs and can trigger expensive enclave paging and frequent world switches. Second, ADS workloads ingest multi-channel, high-bandwidth sensor streams (e.g., camera and LiDAR), so naïvely confining the full perception–planning pipeline to a secure enclave can incur substantial I/O marshalling overhead and may violate real-time latency budgets. Existing TEE-for-AD prototypes are often evaluated on low-bandwidth channels (e.g., GPS) rather than on realistic high-throughput perception inputs (Hu et al., 2020), leaving the end-to-end feasibility for FM-scale workloads under-explored. Third, in today’s deployment reality, widely used commodity accelerators may not expose confidential-computing capabilities suitable for isolating FM execution, while platforms that do provide stronger TEE-style protections can be substantially more costly and harder to provision at scale; this gap motivates practical designs that minimize the trusted computing base (TCB) and carefully trade off security guarantees against latency and integration overhead.
  • Opportunity: Security-by-Design Hardware Architectures. A primary research opportunity lies in architecting hardware with intrinsic security guarantees, with a key approach being the development of Trusted Execution Environments (TEEs) for AI accelerators (Yao et al., 2024a; Chen et al., 2024a). A TEE (Chen et al., 2020) would leverage hardware to create an isolated enclave, protecting the confidentiality and integrity of an FM’s parameters and its execution, even from a compromised high-level operating system. From a deployment perspective, widely used commodity accelerator stacks can provide limited support for enclave-style protection. Prior work reports that commonly used commercial GPUs in current AV/AD solutions/platforms (e.g., RTX A5000 in AVA-3510) do not provide TEE functionality, whereas confidential-computing-enabled accelerators (e.g., Hopper-class GPUs) can be substantially more costly (Li et al., 2025c). This deployment gap motivates hybrid designs that minimize the trusted computing base (TCB) and confine only security-critical components to the trusted domain. Concretely, TEEs for FM-powered ADSs must satisfy architectural requirements such as (i) low overhead and real-time predictability (bounded enclave transitions and secure I/O costs), (ii) scalability to FM-scale working sets (or secure partitioning/streaming of model components), (iii) accelerator-aware trust boundaries that account for GPU/NPU driver stacks and DMA paths that often remain outside the enclave, and (iv) multi-sensor support for securely ingesting and attesting to high-throughput perception data. Significant research is therefore needed to design novel, low-overhead TEEs that can scale to serve FMs, establishing a verifiable hardware root of trust for critical AI computations in autonomous driving. To make the comparison space explicit, as shown in Table 4, TEEs can be classified along two orthogonal axes—software- vs. hardware-based and privileged vs. non-privileged deployments—each implying different isolation granularity, attestation support, and performance costs (Muñoz et al., 2023). A systematic evaluation of these design points for FM workloads (secure memory limits, sensor I/O throughput, and accelerator integration) remains an open problem and is essential for identifying feasible, automotive-grade TEE configurations.
  • Opportunity: Open and Verifiable Hardware Stacks. To address the ”black-box” nature of proprietary hardware, there is an opportunity to develop and promote open-source, verifiable hardware designs or standards for FM acceleration. An open and transparent hardware stack would allow for community-driven security audits, reducing the risk of hidden backdoors or design flaws (Wang et al., 2024b). In addition to openness, an important opportunity is to standardize measurement and attestation interfaces (e.g., for firmware provenance, model binaries, and critical runtime configurations) so that safety cases can incorporate verifiable evidence about the hardware and software supply chain. Such interfaces would enable reproducible security evaluation across vendors and help bridge the gap between research prototypes and deployable, certifiable FM acceleration stacks.
Table 4. TEE implementation categories, characteristics, and representative examples (adapted from Muñoz et al. (Muñoz et al., 2023)).
TEE Category Description and Characteristics Examples
1) Hardware-based & Privileged Description: Rely on dedicated hardware isolation (e.g., ARM TrustZone) and run with high system privileges.
Characteristics: (i) Access to broad system resources; (ii) typically uses a secure monitor; (iii) two domains: Secure World (SW) and Normal World (NW).
Commercial: Qualcomm QSEE; Trustonic t-base; Samsung TZ-RKP; Google Trusty.
Open/Academic: Linaro OPTEE; Microsoft TLR; SafeG; Kinibi_M.
2) Hardware-based & Non-privileged Description: Utilize hardware support (e.g., Intel SGX or AMD SEV) without granting the TEE instance total control over the system.
Characteristics: (i) Supports multiple deployments (new instances can be added without extending the Trusted Computing Base); (ii) often referred to as enclaves.
Commercial: Intel SGX; AMD SEV; TrustICE.
Open/Academic: Sanctum; SecureBlue; Haven; SCONE; Graphene-SGX.
3) Software-based & Privileged Description: Software-enforced TEEs (without dedicated secure hardware elements) that run with high privileges to enforce isolation.
Characteristics: (i) Often relies on hypervisor or kernel-level modifications; (ii) isolation is logical rather than physical.
Open/Academic: Nested Kernel; OpenTEE; MicroTEE; SoftTEE; TrustShadow; SKEE.
4) Software-based & Non-privileged Description: Software-only implementations without elevated system-wide privileges.
Characteristics: (i) Often rely on virtualization layers or application-level sandboxing; (ii) isolate specific processes/data without full system control.
Open/Academic: Overshadow; Virtual Ghost; InkTag; Flicker; TrustVisor; Multizone; Utango.
Challenge II: Resource-Aware Engineering. Training and deploying FMs for autonomous driving tasks demand substantial computational resources, often necessitating large clusters of GPUs or TPUs (Hoffmann et al., 2022; Ran et al., 2025). This challenge is significantly exacerbated during in-vehicle deployment, where models must operate efficiently across diverse and resource-constrained hardware platforms. This is a fundamental infrastructure challenge for researchers and developers, opening up the following opportunities:
  • Opportunity: Distributed and Collaborative Training. One key opportunity lies in the development of distributed and collaborative training frameworks (Lin et al., 2018; Moritz et al., 2018). These frameworks could enable multiple smaller computing entities to pool their computational resources, meeting the demands of training large-scale FMs. This approach aims to overcome the barrier of high-cost infrastructure but also promote a more diverse and inclusive development ecosystem (Ran et al., 2025).
  • Opportunity: Resource-Aware Model Search for Foundation Models. Given the diverse and resource-constrained hardware present in autonomous vehicles, resource-aware model search might be a potential direction for efficient deployment. This approach, often leveraging techniques such as neural architecture search (NAS) (Elsken et al., 2019), aims to automatically discover specialized model architectures that optimally balance performance with certain constraints like latency, memory footprint (Gao et al., 2025a).
@@ 3.3. Development and Deployment of FM4AD
Challenge I: Understanding the FM Development Toolchain. The development toolchain for Foundation Models presents unprecedented complexity compared to traditional deep learning frameworks. The enormous scale of these models significantly amplifies the intricacy of data pre-processing pipelines, distributed training systems, and model deployment workflows (Wang et al., 2024b). Compounding this issue is the rapid pace of innovation in the field, which hinders developers and researchers from maintaining a comprehensive understanding of the continuously evolving ecosystem of tools and libraries (Wang et al., 2024b). This complexity and opacity create significant opportunities to systematically analyze and improve how FMs are built and maintained.
  • Opportunity: Empirical Analysis for Toolchain Optimization. A key research opportunity lies in the large-scale empirical analysis of the FM development toolchain. By systematically examining public repositories, development workflows, and the usage of popular libraries, researchers can identify common inefficiencies, performance bottlenecks, and resource-intensive anti-patterns (Wang et al., 2024b; Nguyen et al., 2019; Liu et al., 2023a). The insights gained from such studies can lead to data-driven best practices and automated tools that help developers streamline complex processes, optimize resource consumption, and accelerate the overall development lifecycle.
  • Opportunity: Quality Assurance for the FM Development Toolchain. While traditional deep learning frameworks benefit from a mature suite of quality assurance and testing techniques (Xiao et al., 2022; Wei et al., 2022; Zhang et al., 2025b), these methods are often inadequate for the complex, distributed nature of the FM toolchain. Core components for large-scale training, such as Ray, introduce immense challenges related to network failures, asynchronous operations, and state management that conventional debuggers and testing methods cannot easily handle. Although initial research has begun to address these challenges (Lu et al., 2025; Wang et al., 2025a), these efforts are still in their early stages, leaving a critical opportunity to develop a new generation of QA tools and methodologies specifically for FM engineering. Research in this area could focus on scalable debuggers for distributed systems, fault injection frameworks to test resilience, and novel validation techniques for massive data-parallel pipelines, ultimately enhancing the reliability and reproducibility of the entire FM development process.
Challenge II: Efficient and Reliable FM Deployment in Vehicle. With growing concerns over data privacy and the strict low-latency requirements of autonomous driving, the on-board deployment of FMs in vehicles is becoming essential. A significant body of research has focused on making this feasible through model compression techniques like pruning (Ma et al., 2023; Sun et al., 2024a), knowledge distillation (Gu et al., 2024; Kim et al., 2024a), and quantization (Lin et al., 2024; Xi et al., 2025; Li et al., 2025a), as well as inference optimizations such as parallel computation (Shoeybi et al., 2020; Yu et al., 2024) and KV cache management (Kwon et al., 2023; Luohe et al., 2024). However, while these techniques improve efficiency, they can inadvertently introduce new security and reliability vulnerabilities. For instance, researchers have demonstrated tailored attacks that specifically exploit the characteristics of quantized models (Zhang et al., 2025a; Yang et al., 2024b) and KV cache optimizations (Song et al., 2025a; Wu et al., 2025b). This inherent tension between performance and security creates an urgent need for deployment strategies that are both highly efficient and fundamentally trustworthy. Furthermore, to navigate this tension and help practitioners better trade-off in practice, we provide the risk-utility guidelines (shown in Table 5).
Table 5. Risk-Utility assessment framework for FM deployment optimization techniques
Technique Benefit Introduced Risks Mitigations
Quantization Maximize speed and storage reduction; hardware compatibility; no re-training (generally); Increases susceptibility to attacks targeting quantized models (Zhang et al., 2025a; Yang et al., 2024b; Egashira et al., 2024); increases the chance of hallucination (Li et al., 2024a) Safety patching (Chen et al., 2025c); safety-aware quantization-aware training;
Pruning Reduction in computation and memory usage; less risk of hallucination (Chrysostomou et al., 2024) Pruning-activated attacks (Egashira et al., 2026); privacy leakage (Shang et al., 2026; Kuang et al., 2025) Security-aware calibration; model patching with repaired parameters (Egashira et al., 2026); iterative compensation (Frantar and Alistarh, 2023)
Model distillation Flexible student model architecture; high performance retention; Exploitation of imperfections; teacher hacking (Tiapkin et al., 2025) (transfer of unsafe behaviors, potential lack of generalization); Prioritize online data generation (dynamic sampling) and high prompt diversity; limiting re-training epochs;
KV Cache Optimization Enhances real-time responsiveness and throughput; Potential verbatim input reconstruction and semantic exfiltration through side channel attack (Wu et al., 2025b); direct inversion/collision attack (Luo et al., 2025) Obfuscation schemes like KV-Cloak (Luo et al., 2025); execution isolation within TEEs to prevent leakage;
  • Opportunity: Empirical Analysis on the Impact of Optimization Techniques. While specific vulnerabilities in optimized models have been identified (Zhang et al., 2025a; Yang et al., 2024b; Song et al., 2025a; Wu et al., 2025b), there is currently a lack of large-scale empirical studies that systematically measure how these techniques impact a model’s overall trustworthiness in autonomous driving tasks. This creates a vital research opportunity to conduct comprehensive analyses that quantify the effects of pruning, quantization, and other optimizations on key properties beyond performance. Such studies should evaluate the trade-offs concerning adversarial robustness, fairness, reliability on out-of-distribution data, and the model’s propensity for hallucination. The findings would establish an evidence-based understanding for practitioners, leading to clear guidelines for safely applying optimization techniques in safety-critical systems like autonomous vehicles.
  • Opportunity: Robust-by-Design Optimization Techniques. A key research opportunity lies in co-designing model optimization techniques with security and robustness as first-class objectives. Instead of optimizing for performance alone, future research could focus on developing new pruning, distillation, or quantization algorithms that are inherently resistant to known attack vectors. For example, this could involve creating quantization schemes that provably maintain adversarial robustness or knowledge distillation processes that transfer security properties from a larger teacher model to a smaller student model. The goal is to create a new class of optimization methods where efficiency gains do not come at the cost of safety.
Challenge III: Model Maintenance and Lifecycle Management. The rapid evolution and high computational cost of FMs create a critical need for structured maintenance and lifecycle management within the ADS development process. A major engineering hurdle is the resource-intensive nature of model training; it is often computationally prohibitive to retrain large-scale models from scratch to adapt to new environmental shifts or specific task variations. Current practices frequently lack systematic frameworks for model versioning, discovery, and reuse. This results in significant redundant effort and inefficiency within the FM supply chain (Wang et al., 2024b), as developers may lack the tools to identify and retrieve existing pre-trained models with similar capabilities that could be adapted through modular updates or lightweight fine-tuning rather than full retraining (Ren et al., 2023).
  • Opportunity: Model Recommendation and Reuse Frameworks. To mitigate the maintenance overhead, there is a vital research opportunity in developing automated model recommendation engines for the FM-based ADS ecosystem. These frameworks could utilize metadata and gradient-based fingerprinting techniques, such as TENSORGUARD (Wu et al., 2025c), to help developers discover similar FMs within a curated repository that best match a target domain’s requirements. By treating FMs as software artifacts requiring provenance tracking, researchers can enable similarity detection and family classification independently of training data or specific model formats (Wu et al., 2025c).
Challenge IV: Edge/Cloud Collaboration for FM Services. Deploying FMs in autonomous vehicles presents a fundamental trade-off between on-board (edge) and remote (cloud) computations. While optimization techniques like quantization can reduce the burden on edge devices, their limited computing power inherently caps model capability and complexity. Conversely, cloud infrastructure offers vast computational resources for complex reasoning but cannot meet the strict low-latency, high-reliability, and data privacy requirements essential for safety-critical driving decisions. This gap, where neither edge nor cloud alone provides a complete solution, creates a significant opportunity to design hybrid systems that strategically leverage the strengths of both environments.
  • Opportunity: Intelligent Task Orchestration. A primary opportunity lies in creating intelligent frameworks that dynamically schedule and offload tasks across the edge-cloud continuum (Yang et al., 2024c; Hao et al., 2024). Research in this area focuses on developing algorithms that can partition workloads in real-time: latency-critical functions like immediate hazard detection would remain on the edge, while computationally intensive, non-real-time tasks like complex scene interpretation or HD map updates could be sent to the cloud. The goal is to build an adaptive system that optimizes for performance, latency, and resource utilization based on current driving context and network conditions.
  • Opportunity: Robust and Asynchronous Communication. The connection between a vehicle and the cloud is often intermittent and variable. A key opportunity is to design robust and asynchronous communication schemes that ensure system reliability despite unstable network connectivity (Zhang et al., 2024a). This includes developing mechanisms that allow the edge model to operate autonomously when disconnected and then asynchronously sync its knowledge or receive updates from the cloud when a connection is re-established. Such schemes are vital for ensuring the vehicle remains safe and operational at all times while still benefiting from the power of the cloud.